On February 1, 2018, Sanford & Son issued 12% bonds dated February 1, 2018, with a face amount of $100,000. The bonds sold for $117,160 and mature in 20 years. The effective interest rate for these bonds was 10%. Interest is paid semiannually on July 31 and January 31. Sanford & Son's fiscal year is the calendar year.

Required:
1. Prepare the journal entry to record the bond issuance on February 1, 2018.
2. Prepare the entry to record interest on July 31, 2018, using the straight-line method.
3. Prepare the necessary journal entry on December 31, 2018.
4. Prepare the necessary journal entry on January 31, 2019.
Answer #1
1 February 1, 2018 Cash 117160
      Premium on bonds payable 17160
      Bonds payable 100000
2 July 31, 2018 Interest expense 5571
Premium on bonds payable 429 =17160/40
     Cash 6000 =100000*12%/2
3 December 31, 2018. Interest expense 4643
Premium on bonds payable 358 =429*5/6
     Interest payable 5000 =100000*12%*5/12
4 January 31, 2019 Interest expense 929
Interest payable 5000
Premium on bonds payable 71 =429-358
       Cash 6000 =100000*12%/2
Note: Answers might vary +1 due to rounding off